# A mathematical theory of topological invariants of quantum spin systems

### v2

@@ 6.2 Construction of topological invariants
It suffices to consider the case $A={\mathbb{R}}$. Let ${\mathsf{F}}\in{\mathfrak{D}}^{\psi}_{al}(U)$ and ${\mathsf{G}}\in{\mathfrak{D}}_{al}(V)$. Propositions 3.2 and 3.6 give
$\psi([{\mathsf{F}},{\mathsf{G}}])=\sum_{X\cap Y\neq\varnothing}\psi([{\mathsf{F}}^{X},{\mathsf{G}}^{Y}])=\sum_{Y}\psi({\mathsf{F}}({\mathsf{G}}^{Y}))=0,$
where the last equality is because ${\mathsf{F}}$ preserves $\psi$. ∎
In what follows we will apply Lemma 6.1 to the graded local Lie algebra ${\mathfrak{D}}^{\psi}_{al}\otimes A$ (see Remark 6.1). The evaluation of cycles of $C^{aug}_{\bullet}({\mathfrak{U}},W;{\mathfrak{D}}^{\psi}_{al}\otimes A)$ on spherical CS cocycles has especially nice properties when DGLA cycles belong to the commutator DGLA
$\overline{\left[C^{aug}_{\bullet+1}\left({\mathfrak{U}},W;{\mathfrak{D}}^{\psi}_{al}\otimes A\right),C^{aug}_{\bullet+1}\left({\mathfrak{U}},W,{\mathfrak{D}}^{\psi}_{al}\otimes A\right)\right]}.$ (24)
Proposition 6.3.
If a cycle ${\mathsf{f}}$ lies in the commutator DGLA (24), then $\psi(\langle{\mathsf{f}},\beta\rangle)$ depends only on the cohomology class of $\beta$ and the homology class of ${\mathsf{f}}$ in the DGLA (24).
@@
Suppose ${\mathsf{f}}\in C^{aug}_{p}({\mathfrak{U}},W;{\mathfrak{D}}^{\psi}_{al}\otimes A)$ and ${\mathsf{f}}=\partial{\mathsf{g}}$ for some ${\mathsf{g}}$ in the commutator DGLA. Then the vanishing of $\psi(\langle{\mathsf{f}},\beta\rangle)\in A$ follows from the proof of Lemma 6.1 and Lemma 6.2. Thus it remains to show that $\psi(\langle{\mathsf{f}},\check{\delta}b\rangle)=0$ for any $b\in C^{p-1}({\mathfrak{U}},\hat{W})$ if ${\mathsf{f}}$ is a cycle of the commutator DGLA. Indeed, since $\langle{\mathsf{f}},\partial^{*}b\rangle=\langle\partial{\mathsf{f}},b\rangle=0$, we have
@@
$\langle{\mathsf{f}},\check{\delta}b\rangle=\sum_{\begin{subarray}{c}s\in I_{p},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{unbounded}\end{subarray}}{\mathsf{f}}_{s}(\partial^{*}b)_{s}=-\sum_{\begin{subarray}{c}s\in I_{p},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{bounded}\end{subarray}}{\mathsf{f}}_{s}(\partial^{*}b)_{s},$
and the $\psi$-average of each term in the latter sum vanishes by Lemma 6.2. ∎
The evaluations of cycles of the commutator DGLA on spherical CS cocycles can be used to construct topological invariants of $G$-invariant gapped states on ${\mathbb{R}}^{n}$. We set $W={\mathbb{R}}^{n}$. The cycle we use as an input is the commutator class of $C_{\bullet}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al})$ defined using the inhomogeneous Maurer-Cartan equation (Section 6.1). Since ${\mathfrak{D}}^{\psi,\bf G}_{al}$ is a sub-algebra of ${\mathfrak{D}}^{\psi}_{al}\otimes A$ with $A=\prod_{k=1}^{\infty}{\rm Sym}^{k}{\mathfrak{g}}^{*}[-2]$, Prop. 6.3 applies to such cycles.
Let ${\mathfrak{U}}$ be a CS cover of ${\mathbb{R}}^{n}$, and $\beta$ be a spherical CS cocycle of $\hat{W}=S^{n-1}$ of degree $l$ with respect to the cover $\hat{\mathfrak{U}}$. Let ${\mathsf{p}}$ be a ${\mathsf{Q}}$-twisted MC element of $C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al})$. Evaluating $[{\mathsf{p}},{\mathsf{p}}]$ on $\beta$ and taking into account that the grading is shifted by $1$ relative to the Čech grading, we get a $G$-invariant element of ${{\mathfrak{d}}^{\psi}_{al}}\otimes{\rm Sym}^{k}({\mathfrak{g}}^{*}[-2])$ where $2k-3=l$. Since $\psi$ is $G$-invariant, $\psi(\langle[{\mathsf{p}},{\mathsf{p}}],\beta\rangle)$ is a $G$-invariant element of ${\rm Sym}^{k}({\mathfrak{g}}^{*}[-2])$. Equivalently, it is the value of a degree $3$ linear function on cocycles of $\check{C}^{\bullet}(\hat{\mathfrak{U}},{\mathbb{R}})$ valued in ${\rm Sym}^{\bullet}{\mathfrak{g}}^{*}[2]$.
Theorem 6.1.
The function $\beta\mapsto\psi(\langle[{\mathsf{p}},{\mathsf{p}}],\beta\rangle)$ depends only on $\psi$ and the class of $({\mathfrak{U}},\beta)$ in the CS cohomology of $S^{n-1}$.
@@
For a fixed covering ${\mathfrak{U}}$, Prop. 6.1 and 6.3 implies $\psi(\langle[{\mathsf{p}},{\mathsf{p}}],\beta\rangle)$ is independent of ${\mathsf{p}}$. It depends on $\beta$ solely through its cohomology class. It remains to show invariance under refinement of cover.
Let $({\mathfrak{V}},\phi)$ be a refinement of ${\mathfrak{U}}$ with $V_{j}\subset U_{\phi(j)}.$ Cocycles $({\mathfrak{U}},\beta)$ and $({\mathfrak{V}},\phi^{*}\beta)$, where $(\phi^{*}\beta)_{j_{0},\dots,j_{k}}=\beta_{\phi(j_{0}),\dots,\phi(j_{k})}$, are in the same CS cohomology class. From Prop.6.1 there exists ${\mathsf{Q}}$-twisted MC element ${\mathsf{p}}$ for the cover ${\mathfrak{V}}$. Prop.6.2 then implies that $\phi_{*}{\mathsf{p}}$ is a ${\mathsf{Q}}$-twisted MC element ${\mathsf{p}}$ for the cover ${\mathfrak{U}}$. An easy expansion shows
$\langle[\phi_{*}{\mathsf{p}},\phi_{*}{\mathsf{p}}],\beta\rangle=\langle[{\mathsf{p}},{\mathsf{p}}],\phi^{*}\beta\rangle.$
By Proposition 4.7, the cohomology group $H^{l}_{CS}(S^{n-1},{\mathbb{R}})$ is one-dimensioal for $l=n-1$ and vanishes otherwise. It is natural to take the class of $({\mathfrak{U}},\beta)$ to be a generator of $H^{n-1}_{CS}(S^{n-1},{\mathbb{Z}})$. This generator is uniquely defined once an orientation of ${\mathbb{R}}^{n-1}$ has been chosen. Thus for any even $n$ we obtain an invariant of gapped $G$-invariant states on ${\mathbb{R}}^{n}$ taking values in $G$-invariant polynomials on ${\mathfrak{g}}$ of degree $(n+2)/2$, and this invariant changes sign when the orientation of ${\mathbb{R}}^{n}$ is changed. This is in agreement with Chern-Simons field theory.
@@ 6.3 The Hall conductance
Let $\psi$ be a $U(1)$-invariant gapped state of a lattice system on ${\mathbb{R}}^{n}$. The generator of the $U(1)$ action is ${\mathsf{Q}}\in{\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$. Let ${\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}$ be the sub-algebra of $U(1)$-invariant elements of ${\mathfrak{D}}^{\psi}_{al}$. Since $U(1)$ is connected, this the same as the sub-algebra of elements of ${\mathfrak{D}}_{al}$ which commute with ${\mathsf{Q}}$. More generally, for any $U\in{\mathcal{C}\mathcal{S}}_{n}$ ${\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}(U)={\mathfrak{D}}_{al}(U)\cap{\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}$. This is a local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$. The graded local Lie algebra ${\mathfrak{D}}^{\psi,\bf G}_{al}$ reduces in this case to ${\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}\otimes{\mathbb{R}}[[t]]$ where $t$ is a variable of degree $-2$.
@@
Pick a cover ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ of ${\mathbb{R}}^{n}$. To find a solution ${\mathsf{p}}$ of the inhomogenenous Maurer-Cartan equation with ${\mathsf{B}}={\mathsf{Q}}\otimes t$, we write ${\mathsf{p}}=\sum_{k=1}^{\infty}{\mathsf{p}}_{k}\otimes t^{k}$, where ${\mathsf{p}}_{k}\in C_{k-1}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al})$. To compute the topological invariant of a state on ${\mathbb{R}}^{2}$ it is sufficient to solve for ${\mathsf{p}}_{1}$.
@@
${\mathsf{p}}_{1}$ is a solution of $\partial{\mathsf{p}}_{1}={\mathsf{Q}}$. Explicitly, ${\mathsf{p}}_{1}=\{{\mathsf{Q}}_{i}\in{\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}(U_{i})\}_{i\in I}$ such that $\sum_{i\in I}{\mathsf{Q}}_{i}={\mathsf{Q}}$. Such ${\mathsf{Q}}_{i}$ exist because ${\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}$ is a cosheaf. Then the component of the commutator class in $C_{1}(U,{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al})$ is $\{[{\mathsf{Q}}_{i},{\mathsf{Q}}_{j}]\}_{i<j}$. The topological invariant of the state $\psi$ is obtained by evaluating it on a Čech 1-cocycle $\beta$ on $S^{1}$ corresponding to the cover $\hat{\mathfrak{U}}$ and then averaging the resulting inner derivation:
$\sigma(\beta)=\psi\left(\sum_{i<j}\beta_{ij}[{\mathsf{Q}}_{i},{\mathsf{Q}}_{j}]\right).$
@@
Note that averaging over $\psi$ must be performed after the summation over $i,j$ because $[{\mathsf{Q}}_{i},{\mathsf{Q}}_{j}]$ is not an inner derivation, in general.
The simplest CS cover of ${\mathbb{R}}^{2}$ which can represent a nontrivial class in $H^{1}_{CS}(S^{1},{\mathbb{R}})$ is made of three cones with a common vertex. In this case the construction of the invariant reduces to that in [7, 11]. It was shown there that the resulting invariant is proportional to the zero-temperature Hall conductance as determined by the Kubo formula.

### v3

@@ 6.2 Construction of topological invariants
It suffices to consider the case $A={\mathbb{R}}$. We begin by showing that the sum
$\sum_{X,Y,Z\in{\mathbb{B}}_{n}}\pi^{Z}([{\mathsf{F}}^{X},{\mathsf{G}}^{Y}])$ (134)
is absolutely convergent. Using Lemmas 11 and 31
$\sum_{X,Y\in{\mathbb{B}}_{n}}\sum_{Z\in{\mathbb{B}}_{n}}\|\pi^{Z}([{\mathsf{F}}^{X},{\mathsf{G}}^{Y}])\|$ $\leq\sum_{X,Y\in{\mathbb{B}}_{n}}\sum_{Z\subseteq X\wedge Y}4^{d}2\|{\mathsf{F}}^{X}\|\|{\mathsf{G}}^{Y}\|$
$\leq\sum_{X,Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(X\wedge Y))^{2d}4^{d}2\|{\mathsf{F}}^{X}\|\|{\mathsf{G}}^{Y}\|$
$\leq\sum_{X,Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(X)+\operatorname{diam}(Y))^{2d}4^{d}2\|{\mathsf{F}}^{X}\|\|{\mathsf{G}}^{Y}\|$
$\leq 4^{d}2\left(\sum_{X\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(X))^{2d}\|{\mathsf{F}}^{X}\|\right)\left(\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y))^{2d}\|{\mathsf{G}}^{Y}\|\right)$
$<\infty$ (135)
Recall that for an anchored derivation $\mathsf{H}$} we define $\psi({\mathsf{H}})=\psi(\Sigma({\mathsf{H}}))$. We have
$\psi(\Sigma([{\mathsf{F}},{\mathsf{G}}]))$ $=\sum_{X,Y,Z\in{\mathbb{B}}_{n}}\psi(\pi^{Z}([{\mathsf{F}}^{X},{\mathsf{G}}^{Y}])$
$=\sum_{X,Y\in{\mathbb{B}}_{n}}\psi([{\mathsf{F}}^{X},{\mathsf{G}}^{Y}])$
$=\sum_{Y\in{\mathbb{B}}_{n}}\psi({\mathsf{F}}({\mathsf{G}}^{Y}))$
$=0.$ (136)
Lemma 114.
[…]
@@
[…]
@@
$\langle{\mathsf{f}},\check{\delta}b\rangle=\sum_{\begin{subarray}{c}s\in I^{p+1},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{unbounded}\end{subarray}}{\mathsf{f}}_{s}(\partial^{*}b)_{s}=-\sum_{\begin{subarray}{c}s\in I^{p+1},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{bounded}\end{subarray}}{\mathsf{f}}_{s}(\partial^{*}b)_{s},$ (137)
and the $\psi$-average of each term in the latter sum vanishes by Lemma 113. ∎
Proof of Theorem 112.
Note first that the anchored derivation ${\mathsf{p}},{\mathsf{p}}]$ is anti-hermitian, hence its expectation value is purely imaginary, and $\nu_{\psi}(\beta)$ is real. Also, since the commutator class has even degree and thus odd Čech degree, $\nu_{\psi}(\beta)=0$ if $\beta$ has even degree. For a fixed covering $\mathfrak{U}$}, Prop. 105 and Lemma 114 imply $\psi(\Sigma\langle[{\mathsf{p}},{\mathsf{p}}],\beta\rangle)$ is independent of the choice of the solution $\mathsf{p}$} of the twisted MC equation. By Lemma 114, it depends on the cocycle $\beta$ solely through its cohomology class. It remains to show invariance under refinement of the covering $\mathfrak{U}$}.
Let $({\mathfrak{V}},\phi)$ be a refinement of $\mathfrak{U}$} with $V_{j}\leq U_{\phi(j)}.$ Cocycles $({\mathfrak{U}},\beta)$ and $({\mathfrak{V}},\phi^{*}\beta)$, where $(\phi^{*}\beta)_{j_{0},\dots,j_{k}}=\beta_{\phi(j_{0}),\dots,\phi(j_{k})}$, are in the same CS cohomology class. From Prop.105 there exists $\mathsf{Q}$}-twisted MC element $\mathsf{p}$} for the cover $\mathfrak{V}$}. Prop.106 then implies that $\phi_{*}{\mathsf{p}}$ is a $\mathsf{Q}$}-twisted MC element for the cover $\mathfrak{U}$}. An easy expansion shows
$\langle[\phi_{*}{\mathsf{p}},\phi_{*}{\mathsf{p}}],\beta\rangle=\langle[{\mathsf{p}},{\mathsf{p}}],\phi^{*}\beta\rangle.$ (138)
[…]
In the case $W={\mathbb{R}}^{n}$, the odd cohomology is nontrivial only for even $n$ and is a one-dimensional vector space of degree n1n-1. Once the orientation of ${\mathbb{R}}^{n}$ has been chosen, there is a canonical basis element, the dual of the fundamental class of $S^{n-1}$. Thus, for $G$-invariant gapped states of quantum lattice systems on ${\mathbb{R}}^{n}$ with $n$ even, Theorem 112 supplies a unique invariant taking values in $G$-invariant polynomials on $\mathfrak{g}$} of degree $(n+2)/2$. This invariant changes sign when the orientation of ${\mathbb{R}}^{n}$ is changed. Note that the same datum classifies characteristic classes of $G$-bundles with real coefficients in degree $n+2$, or equivalently, classical Chern-Simons field theories with gauge group $G$ in space-time dimension $n+1$.
In Section 2 we defined the notion of a smooth path of gapped states. One can extend it to $G$-invariant gapped states in an obvious manner.
Definition 115.
Fix an on-site action of a compact Lie group $G$ on a quantum lattice system. A smooth path of $G$-invariant LGAs is a smooth path of LGAs $u\mapsto\alpha(u)$ such that $\alpha(u)$ commutes with the $G$-action for all $u\in[0,1]$. A smooth path of $G$-invariant gapped states is a smooth path of states of the form $u\mapsto\psi(u)=\omega\circ\alpha(u)$, where $\omega$ is a $G$-invariant gapped state and $u\mapsto\alpha(u)$ is a smooth path of $G$-invariant LGAs.
The following proposition shows that the function $\nu_{\psi}$ does not vary in smooth families of states. This justifies its interpretation as a topological invariant of the state $\psi$.
Proposition 116.
Let $u\mapsto\psi(u)$ be a smooth path of $G$-invariant gapped states of a quantum lattice system on $W\in{\mathcal{C}\mathcal{S}}_{n}$. Then $\nu_{\psi(u)}=\nu_{\psi(0)}$ for all $u\in[0,1]$.
@@
By definition of a smooth path, for any $u\in[0,1]$ there exists a smooth path of LGAs $u\mapsto\alpha(u)$ commuting with the $G$-action such that $\psi(u)=\psi(0)\circ\alpha(u)$. Therefore, ${\mathfrak{D}}_{al}^{\psi(u),{\mathbf{G}}}(U)=\alpha(u)^{-1}({\mathfrak{D}}_{al}^{\psi(0),{\mathbf{G}}}(U))$ for every $U\in{\mathcal{C}\mathcal{S}}_{n}$. Hence for every $\mathcal{C}\mathcal{S}$} cover $\mathfrak{U}$} the graded DGLAs corresponding to these two local Lie systems are related by conjugation with $\alpha$ inside $C^{aug}_{\bullet+1}({\mathfrak{U}},W;{\mathfrak{D}}_{al}^{\mathbf{G}})$. Furthermore, these DGLAs are pointed and their curvatures coincide. Indeed, the curvature is the generator $\mathsf{Q}$} of the $G$-action regarded as the distinguished degree 2-2 element of ${\mathfrak{D}}_{al}^{\mathbf{G}}(W)$ which preserves $\psi(u)$ for all $u\in[0,1]$. Finally, if ${\mathsf{p}}(0)$ is a twisted MC element corresponding to $\psi(0)$, then $\alpha(u)^{-1}({\mathsf{p}}(0))$ is a twisted MC element for $\psi(u)$. This implies $\nu_{\psi(u)}=\nu_{\psi(0)}$. ∎
@@ 6.3 The Hall conductance
Let $\psi$ be a $U(1)$-invariant gapped state of a lattice system on ${\mathbb{R}}^{n}$. The generator of the $U(1)$ action is ${\mathsf{Q}}\in{\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$. Let ${\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}$ be the sub-algebra of $U(1)$-invariant elements of ${\mathfrak{D}}^{\psi}_{al}$. Since $U(1)$ is connected, this is the same as the sub-algebra of elements of ${\mathfrak{D}}_{al}$ which commute with $\mathsf{Q}$}. More generally, for any $U\in{\mathcal{C}\mathcal{S}}_{n}$ ${\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}(U)={\mathfrak{D}}_{al}(U)\cap{\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}$. This is a local Lie system over ${\mathcal{C}\mathcal{S}}_{n}$. The graded local Lie system ${\mathfrak{D}}^{\psi,\bf G}_{al}$ reduces in this case to ${\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}\otimes{\mathbb{R}}[[t]]$ where $t$ is a variable of degree 2-2.
@@
Pick a cover ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ of ${\mathbb{R}}^{n}$. To find a solution $\mathsf{p}$} of the inhomogeneous Maurer-Cartan equation with ${\mathsf{B}}={\mathsf{Q}}\otimes t$, we write ${\mathsf{p}}=\sum_{k=1}^{\infty}{\mathsf{p}}_{k}\otimes t^{k}$, where ${\mathsf{p}}_{k}\in C_{k-1}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al})$. To compute the topological invariant of a state on ${\mathbb{R}}^{2}$ it is sufficient to solve for ${\mathsf{p}}_{1}$.
@@
${\mathsf{p}}_{1}$ is a solution of $\partial{\mathsf{p}}_{1}={\mathsf{Q}}$. Explicitly, ${\mathsf{p}}_{1}=\{{\mathsf{Q}}_{i}\in{\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}(U_{i})\}_{i\in I}$ such that $\sum_{i\in I}{\mathsf{Q}}_{i}={\mathsf{Q}}$. Such ${\mathsf{Q}}_{i}$ exist because ${\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al}$ is a cosheaf. Then the component of the commutator class in $C_{1}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,{\mathsf{Q}}}_{al})$ is $\oplus_{i,j}[{\mathsf{Q}}_{i},{\mathsf{Q}}_{j}]$. The topological invariant of the state $\psi$ is obtained by evaluating it on a Čech 1-cocycle $\beta$ on $S^{1}$ corresponding to the cover $\hat{\mathfrak{U}}$ and then averaging the resulting anchored derivation:
$\nu_{\psi}(\beta)={\sqrt{-1}}\psi\left(\sum_{i,j}\beta_{ij}[{\mathsf{Q}}_{i},{\mathsf{Q}}_{j}]\right).$ (139)
@@
Note that averaging over $\psi$ must be performed after the summation over $i,j$ because ${\mathsf{Q}}_{i},{\mathsf{Q}}_{j}]$ is not an anchored derivation, in general.
The simplest cover of ${\mathbb{R}}^{2}$ which can represent a nontrivial class in $H^{1}_{CS}(S^{1},{\mathbb{R}})$ involves three cones $U_{i}$, $i\in\{1,2,3\},$ with a common vertex. A cocycle representing the canonical generator of $H^{1}(S^{1},{\mathbb{Z}})$ is $\beta_{12}=-\beta_{21}=1$, $\beta_{13}=\beta_{23}=0$. Then the invariant is given by
$\nu_{\psi}(\beta)=2{\sqrt{-1}}\psi(\Sigma[{\mathsf{Q}}_{1},{\mathsf{Q}}_{2}])\in{\mathbb{R}}.$ (140)
It was shown in Section 4.1 of [18] (see also Section 4.4.2 of [19]) that $\nu_{\psi}(\beta)$ is the zero-temperature Hall conductance of the system in units of $e^{2}/\hbar$, where $e$ is the electron charge and $\hbar$ is the Planck constant.
Example 117.
Consider a gapped Fock state of the CAR algebra associated with a lattice $\Lambda\subseteq{\mathbb{R}}^{2}$ and a single-site Hilbert space $H$ (Example 48). It depends on a projector $P$ acting on $\ell^{2}(\Lambda,H)$ . If the cover $U_{i}$, $i=1,2,3$, is chosen as above, one can take ${\mathsf{Q}}_{i}$ to have the form ${\mathsf{F}}(h_{i})$ where $h_{i}$ are operators on $\ell^{2}(\Lambda,H)$ given by (98). Evaluating the expectation value (140), one finds
$-{\sqrt{-1}}\nu_{\psi}(\beta)=2{\rm Tr}\,P[h_{1},h_{2}]=2{\rm Tr}\,(Pf_{1}Pf_{2}P-Pf_{2}Pf_{1}P)=2{\rm Tr}[Pf_{1}P,Pf_{2}P],$ (141)
where the functions $f_{i}:\Lambda\rightarrow{\mathbb{R}}$ are given by (97). Note that the operators $Pf_{i}P$ are not trace class, so one cannot use the cyclic property of the trace to conclude that $\nu_{\psi}(\beta)$ is zero. It is shown in Appendix C.3 of [21] that for any gapped Fock state $2\pi\nu_{\psi}(\beta)$ is an integer. Examples are known where this integer is nonzero, see e.g. [36].
Remark 118.
It is a long-standing conjecture that for an arbitrary $U(1)$-invariant gapped state of a quantum lattice system the topological invariant $2\pi\nu_{\psi}(\beta)$ is a rational number.